# LLMs Think, But Not In Your Flow: Reasoning-Level Personalization for Black-Box Large Language Models

### v3

@@ Appendix B LLM-based Clustering Details
Iterative PAS cycle ($\leq R_{\max}$ rounds, $R_{\max}=3$). All LLM-based operations in PAS—including proposing candidate factors, assigning features, and handling residuals—are conducted using the same backbone model, gpt-4o-mini [41].
@@
Balancing factor representativeness against the typical feature count per user, we set $L=16$ candidate factors per round and $P_{\max}=8$ as the maximum factors selected in each iteration.
@@
Further implementation details are available in the provided code and in the original paper [34].
Appendix C Experimental Details
C.1 Dataset and Task Details
We conduct evaluations on four personalization tasks: classification, regression, generation, and question answering, sourced from the LaMP benchmark [18] and GlobalOpinionQA (GOQA) [38], each of which presents unique challenges for modeling user-specific decision-making patterns.
@@ C.2 Baselines
  • Retrieval-Augmented Generation (RAG) [18]: Similar to ICL, but the examples are selected using semantic similarity via Contriever, enabling more relevant context injection.
  • Profile-Augmented Generation (PAG[19]: User histories are summarized into natural-language profile descriptions. We use top-10 retrieved histories for summary generation.
  • HYDRA [17]: A plug-and-play framework that uses a reranker module to reorder retrieved in-context examples and an adapter module to select the most suitable response from multiple LLM generations.
  • Fermi [16]: A prompt refinement method that iteratively updates the user prompt using feedback from prior misaligned generations, optimizing input construction over time.
C.3 Implementation Details
@@
The black-box language model that serves as the backbone across all experiments is gpt-4o-mini (gpt-4o-mini-2024-07-18), accessed via the OpenAI API using the LangChain framework [^1] https://www.langchain.com/. All components of RPM—including feature extraction, factor construction, and reasoning generation—are implemented using dedicated prompt templates designed for each subtask, executed through the API. To ensure deterministic outputs and reproducibility, we fix the decoding temperature to 0.0 for all inference steps across methods. This configuration is consistently applied to RPM and all baseline methods during evaluation.
Exceptions are made for baseline methods that explicitly rely on sampling-based generation as part of their original design. HYDRA [17] selects from sampled response candidates using an adapter controller. We follow the original setting and use temperature 1.0 for sampling. Fermi [16] utilizes prompt optimization via feedback-driven sampling. During its prompt search phase, we also apply a temperature of 1.0 to enable diverse candidate generation. These non-deterministic settings are limited strictly to internal sampling stages defined by the original methods. All final outputs for evaluation are generated with temperature 0.0 to ensure consistency across methods. No maximum token limit was enforced, and no truncation-related issues were observed in any instance.
C.4 Human Evaluation
@@
Appendix D Additional Analyses
@@ D.3 Retrieval Strategy
Table 6 confirms that the sample-level concatenation used in RPM consistently outperforms the feature-level variant, which—despite being competitive—often over-weights generic features shared by many samples and fails to leverage how multiple cues jointly characterize the user’s logic. Also, the feature-level approach incurs a cost of $\mathcal{O}\!\bigl{(}N\,|\mathcal{\mathcal{G}}_{q^{\prime}}|\,\overline{|\mathcal{\mathcal{G}}_{q_{i}}|}\bigr{)}$ cosine evaluations ($N$ is the number of stored examples and $\overline{|\mathcal{\mathcal{G}}_{q_{i}}|}$ is the average feature count per example), whereas the sample-level scheme requires only $\mathcal{O}(N)$. Hence the default retrieval is both more effective and far more efficient.
@@
D.4 Cost Analysis
For each user in the GOQA benchmark, we estimate the total API billing charges incurred from factor construction to reasoning-aligned generation using the OpenAI API pricing document, [^3] https://platform.openai.com/docs/pricing and then averaged these estimates to obtain the mean per-user cost. A competitive response-level personalization baseline, HYDRA [17], was excluded because its parameter tuning requires additional GPU resources beyond the API billing model. For Fermi [16], we measured the combined cost of its prompt-optimization loop and subsequent inference. RPM incurred \$0.06 per user, compared to \$0.32 for Fermi, resulting in an 81% lower API cost and demonstrating greater cost-efficiency. Moreover, RPM achieved higher accuracy (85.2 % vs. 80.0 %), demonstrating that building a personalized reasoning-augmented user history once per user is both more cost-efficient and more effective than repeated prompt optimization.
D.5 RPM with Various Black-Box LLMs
@@
RPM on Diverse Black-Box LLMs. To confirm that our framework is not limited to gpt-4o-mini, we conduct the entire pipeline on three additional black-box LLMs: gpt-3.5-turbo (weaker model), gpt-4o (stronger model), and o3-mini (reasoning model) [10, 41]. Experiments were performed on two complementary benchmarks—LaMP-3 (review texts with discrete ratings) and GOQA (multiple-choice survey questions)—so that both label-defined and label-free tasks were covered under a fixed budget. Table 8 (upper block) shows a consistent trend relative to the default backbone: gpt-4o raises scores, gpt-3.5-turbo lowers them modestly, and o3-mini remains close to gpt-4o-mini. Across all backbones, RPM consistently delivers the best overall results.
@@
Appendix E Limitations
Modeling User Decision-Making. Our framework constructs personalized reasoning paths with features and factors, yet we cannot empirically verify that these representations faithfully mirror each user’s true decision process. Rigorous causal inference would require controlled experimental data and access to rich behavioral signals [45, 46, 47]. As a result, we approximate the personal reasoning process via accessible user history and this interpretable representation consistently yields significant personalization performance gains against baselines. We also validated the plausibility and fidelity of the generated reasoning through a dedicated human evaluation study.
Computation Cost. The full pipeline of RPM calls the LLM multiple times, so a non-trivial computational cost is unavoidable. Even when the entire workflow runs on gpt-4o-mini, one of the most affordable commercial models, the resulting personalization performance remains comparable to our default backbone, and the same personalized memory can be re-used by the stronger gpt-4o without reconstruction, demonstrating that strong cross-model transferability (Table 8) offsets the one-time construction cost. A detailed analysis (Appendix D.4) further confirms that RPM incurs far lower API cost than the strongest prompt-optimization baseline while delivering higher quality of personalization.
@@ Appendix E Limitations
Appendix F Case Studies
The following tables provide a step-by-step view of this process. Table 9 presents the features extracted from the query (abstract) in Example 1. Table 10 summarizes the user-level factors structured from previously extracted features in the profile. Table 11 shows the personalized reasoning generated for Example 1, based on the query and gold response (i.e., the original title). Table 12 displays the model-generated reasoning and final response (predicted title) for the given target query. Table 13 compares the reasoning and responses produced by RPM, HYDRA, and Fermi for the same target query. Together, these tables illustrate how the model builds and applies user-specific reasoning paths from raw input to final output, enabling both interpretability and personalization.
Query (Abstract)
Extracted Features
Features from Example 1
Features from Example 2
Features from Example 3
User-Specific Factors with Statistics
Methodology 86 83/86 (96.5%) 176/183 (96.2%) 7/183 (3.8%) 0 (0.0%)
Evaluation 73 56/73 (76.7%) 132/137 (96.4%) 5/137 (3.6%) 0 (0.0%)
Challenges 58 48/58 (82.8%) 68/76 (89.5%) 4/76 (5.3%) 4/76 (5.3%)
Algorithms 35 30/35 (85.7%) 44/47 (93.6%) 3/47 (6.4%) 0 (0.0%)
Performance 65 42/65 (64.6%) 82/88 (93.2%) 5/88 (5.7%) 1/88 (1.1%)
Component Content
Component Content
Example 1 Location privacy: Identifies the core issue addressed.
Privacy-preserving schemes: Frames prior limitations.
Participant-density-aware: Highlights novel contribution.
Multi-pseudonym mechanism: Describes technical solution.
Sybil attacks: Points out a relevant security threat.
Paillier cryptosystem: Technical foundation of the scheme.
Component Content
HYDRA + CoT
Fermi + CoT
RPM (Ours)
@@ Appendix G Prompts
To show the precise instruction prompts we provided to the model on the LaMP-5 benchmark, we present them as follows. Table 14 shows the prompt used to extract all potentially response-influential features from the raw query. Table 15 shows the prompt for proposing factor candidates by clustering the extracted features. Table 16 shows the prompt for assigning each feature to its most semantically relevant factor. Table 17 shows the prompt for evaluating whether each factor influenced the actual response and determining its polarity. Table 18 shows the prompt for building personalized reasoning paths based on features and factors. Table 19 shows the prompt for reasoning-aligned generation with given reasoning examples and target query.

### v4

@@ Appendix C LLM-based Clustering Details
Iterative PAS cycle ($\leq R_{\max}$ rounds, $R_{\max}=3$). All LLM-based operations in PAS—including proposing candidate factors, assigning features, and handling residuals—are conducted using the same backbone model, gpt-4o-mini (Hurst et al., 2024).
@@
The parameter selection for PAS clustering was guided by a principle of prioritizing generalizability and reproducibility over expensive, dataset-specific hyperparameter tuning. This approach was chosen to demonstrate that the strong performance of RPM stems from the core framework itself, not from fine-grained optimization. Following this principle, the standard, validated parameters from the original cited work were adopted. Specifically, to balance factor representativeness against the typical feature count per user, the number of candidate factors per round was set to $L=16$ and the maximum factors selected in each iteration was set to $P_{\max}=8$. This commitment to principled parameter settings ensures the results are both reproducible and provide a robust baseline for future work.
@@
Further implementation details are available in the provided code and in the original paper (Wang et al., 2023).
Appendix D Experimental Details
D.1 Dataset and Task Details
We conduct evaluations on four personalization tasks: classification, regression, generation, and question answering, sourced from the LaMP benchmark (Salemi et al., 2023) and GlobalOpinionQA (GOQA) (Durmus et al., 2023), each of which presents unique challenges for modeling user-specific decision-making patterns.
@@ D.2 Baselines
  • Retrieval-Augmented Generation (RAG) (Salemi et al., 2023): Similar to ICL, but the examples are selected using semantic similarity via Contriever, enabling more relevant context injection.
  • Profile-Augmented Generation (PAG(Richardson et al., 2023): User histories are summarized into natural-language profile descriptions. We use top-10 retrieved histories for summary generation.
  • HYDRA (Zhuang et al., 2024): A plug-and-play framework that uses a reranker module to reorder retrieved in-context examples and an adapter module to select the most suitable response from multiple LLM generations.
  • Fermi (Kim & Yang, 2024): A prompt refinement method that iteratively updates the user prompt using feedback from prior misaligned generations, optimizing input construction over time.
D.3 Implementation Details
@@
The black-box language model that serves as the backbone across all experiments is gpt-4o-mini (gpt-4o-mini-2024-07-18), accessed via the OpenAI API using the LangChain framework [^1] https://www.langchain.com/. All components of RPM—including feature extraction, factor construction, and reasoning generation—are implemented using dedicated prompt templates designed for each subtask, executed through the API. To ensure deterministic outputs and reproducibility, we fix the decoding temperature to 0.0 for all inference steps across methods. However, minor variations in outputs were occasionally observed, likely attributable to the non-deterministic nature of the API provider’s backend services. This configuration is consistently applied to RPM and all baseline methods during evaluation.
Exceptions are made for baseline methods that explicitly rely on sampling-based generation as part of their original design. HYDRA (Zhuang et al., 2024) selects from sampled response candidates using an adapter controller. We follow the original setting and use temperature 1.0 for sampling. Fermi (Kim & Yang, 2024) utilizes prompt optimization via feedback-driven sampling. During its prompt search phase, we also apply a temperature of 1.0 to enable diverse candidate generation. These non-deterministic settings are limited strictly to internal sampling stages defined by the original methods. All final outputs for evaluation are generated with temperature 0.0 to ensure consistency across methods. No maximum token limit was enforced, and no truncation-related issues were observed in any instance.
D.4 Human Evaluation
@@
Appendix E Additional Analyses
@@ E.3 Retrieval Strategy
Table 7 confirms that the sample-level concatenation used in RPM consistently outperforms the feature-level variant, which—despite being competitive—often over-weights generic features shared by many samples and fails to leverage how multiple cues jointly characterize the user’s behavior pattern. Also, the feature-level approach incurs a cost of $\mathcal{O}\!\bigl(N\,|\mathcal{\mathcal{G}}_{q^{\prime}}|\,\overline{|\mathcal{\mathcal{G}}_{q_{i}}|}\bigr)$ cosine evaluations ($N$ is the number of stored examples and $\overline{|\mathcal{\mathcal{G}}_{q_{i}}|}$ is the average feature count per example), whereas the sample-level scheme requires only $\mathcal{O}(N)$. Hence the default retrieval is both more effective and far more efficient.
@@
E.4 Cost Analysis
To directly address concerns about scalability and efficiency, we provide a transparent and comparative analysis of both inference latency and computational cost on the GOQA benchmark. RPM introduces a small increase in per-user inference time compared to lightweight baselines, but this overhead is minimal and brings substantial gains in accuracy and personalization. For the cost analysis, we employ the same model (GPT-4o-mini).
Inference Latency.
ICL/RAG achieves a per-user inference time of 0.04s, while RPM requires 0.10s, consisting of 0.04s for feature extraction and 0.06s for reasoning-aligned generation. This additional $\sim$0.06s is required for structured reasoning, and all LLM calls are processed asynchronously to minimize bottlenecks. Importantly, after the initial LLM call per user query, retrieval and ranking proceed with efficient feature-based search, requiring no further LLM calls.
Computational Overhead.
RPM incurs a one-time preprocessing cost of \$0.058 per user, which is substantially lower than prompt-heavy or parameter-tuning approaches such as Fermi (\$0.32) and HYDRA (\$0.47 + additional GPU training). At inference, the per-instance cost of RPM (\$0.0037) is slightly higher than ICL/RAG, but remains significantly lower than advanced baselines while delivering superior accuracy.
Summary of Trade-off.
Overall, RPM achieves 85.2% accuracy, outperforming all compared methods, while introducing only a minor increase in latency and cost relative to ICL/RAG. At the same time, it remains far more efficient and scalable than high-overhead methods like Fermi and HYDRA, which require repeated prompt optimization or parameter tuning. This demonstrates that building a personalized reasoning-augmented user history once per user provides a practical and effective trade-off between efficiency and accuracy.
Table 9: Token usage and per-user cost on the GOQA benchmark.
Method Preprocessing Cost (\$) Inference Cost (\$) Accuracy
Zero-shot 0 0.0002 0.562
ICL/RAG 0 0.0007 0.695 / 0.773
PAG 0 0.0013 0.795
HYDRA 0.4679 + Param. Training 0.0028 + Reranker/Adapter 0.800
Fermi 0.3204 0.0007 0.800
RPM (ours) 0.0581 0.0037 0.852
E.5 RPM with Various Black-Box LLMs
@@
RPM on Diverse Black-Box LLMs. To confirm that our framework is not limited to gpt-4o-mini, we conduct the entire pipeline on three additional black-box LLMs: gpt-3.5-turbo (weaker model), gpt-4o (stronger model), and o3-mini (reasoning model) (Achiam et al., 2023; Hurst et al., 2024). Experiments were performed on two complementary benchmarks—LaMP-3 (review texts with discrete ratings) and GOQA (multiple-choice survey questions)—so that both label-defined and label-free tasks were covered under a fixed budget. Table 10 (upper block) shows a consistent trend relative to the default backbone: gpt-4o raises scores, gpt-3.5-turbo lowers them modestly, and o3-mini remains close to gpt-4o-mini. Across all backbones, RPM consistently delivers the best overall results.
@@
Appendix F Limitations
Modeling User-Specific Decision Making Patterns. Since a user’s actual internal thought process is not directly observable, the proposed framework aims to construct a pragmatic approximation of it. This effort involves leveraging accessible data—such as behavioral patterns, contextual information, and responses from user history—to build an explicit and interpretable reasoning model. The effectiveness of this approach is demonstrated in two ways: first, the resulting representations consistently yield significant personalization performance gains against baselines. Second, a dedicated human evaluation study validated the plausibility and faithfulness of the generated reasoning.
Computation Cost. The full pipeline of RPM calls the LLM multiple times, so a non-trivial computational cost is unavoidable. Even when the entire workflow runs on gpt-4o-mini, one of the most affordable commercial models, the resulting personalization performance remains comparable to our default backbone, and the same personalized memory can be re-used by the stronger gpt-4o without reconstruction, demonstrating that strong cross-model transferability (Table 10) offsets the one-time construction cost. A detailed analysis (Appendix E.4) further confirms that RPM incurs far lower API cost than the strongest prompt-optimization baseline while delivering higher quality of personalization.
@@ Appendix G Prompts
To show the precise instruction prompts we provided to the model on the LaMP-5 benchmark, we present them as follows. Table 11 presents the core prompt template designed for generalizability. Table 12 shows the prompt used to extract all potentially response-influential features from the raw query. Table 13 shows the prompt for proposing factor candidates by clustering the extracted features. Table 14 shows the prompt for assigning each feature to its most semantically relevant factor. Table 15 shows the prompt for evaluating whether each factor influenced the actual response and determining its polarity. Table 16 shows the prompt for building personalized reasoning paths based on features and factors. Table 17 shows the prompt for reasoning-aligned generation with given reasoning examples and target query.
Table 11: Example Prompt Template.
Exemplars: {reasoning_examples}
You are an expert in personalized [TASK].
Your task is to predict [TASK_OUTPUT] based on their previous preferences and [TASK_INPUT].
TASK_DESCRIPTION: A brief, one-sentence description of the task’s objective. [TASK_DESCRIPTION]
Analyze the person’s preference factors and statistics:
- Identify which factors strongly influence this person’s [TASK_OUTPUT] preferences
- Note the typical patterns associated with each factor
- Consider the person’s historical [TASK_OUTPUT] preferences as your baseline
- Compare this abstract with the similar examples
- Look for patterns in how specific features influenced [TASK] in the past
- Consider how the previous reasoning might apply to this [TASK_INPUT]
Develop your title prediction by:
- Starting with an understanding of the person’s historical [TASK_OUTPUT] style as a baseline
- Applying successful personalized reasoning from exemplars
- Ensuring the [TASK_OUTPUT] accurately reflects the content of the abstract while matching the person’s style
Return as JSON: { "reasoning": "", "predicted_title": "" }
User Preference Factors and Statistics: {factors}
[TASK_INPUT]: {task input}
[TASK_INPUT] Features: {features}
@@ 3. Develop your title prediction by:
Appendix H Case Studies
The following tables provide a step-by-step view of this process. Table 18 presents the features extracted from the query (abstract) in Example 1. Table 19 summarizes the user-level factors structured from previously extracted features in the profile. Table 20 shows the personalized reasoning generated for Example 1, based on the query and gold response (i.e., the original title). Table 21 displays the model-generated reasoning and final response (predicted title) for the given target query. Table H compares the reasoning and responses produced by RPM, HYDRA, and Fermi for the same target query. Together, these tables illustrate how the model builds and applies user-specific reasoning paths from raw input to final output, enabling both interpretability and personalization.
Query (Abstract)
Extracted Features
Features from Example 1
Features from Example 2
Features from Example 3
User-Specific Factors with Statistics
Methodology 86 83/86 (96.5%) 176/183 (96.2%) 7/183 (3.8%) 0 (0.0%)
Evaluation 73 56/73 (76.7%) 132/137 (96.4%) 5/137 (3.6%) 0 (0.0%)
Challenges 58 48/58 (82.8%) 68/76 (89.5%) 4/76 (5.3%) 4/76 (5.3%)
Algorithms 35 30/35 (85.7%) 44/47 (93.6%) 3/47 (6.4%) 0 (0.0%)
Performance 65 42/65 (64.6%) 82/88 (93.2%) 5/88 (5.7%) 1/88 (1.1%)
Component Content
Component Content
Example 1 Location privacy: Identifies the core issue addressed.
Privacy-preserving schemes: Frames prior limitations.
Participant-density-aware: Highlights novel contribution.
Multi-pseudonym mechanism: Describes technical solution.
Sybil attacks: Points out a relevant security threat.
Paillier cryptosystem: Technical foundation of the scheme.
Component Content
[…]